# Understanding Catastrophic Forgetting in Language Models via Implicit Inference

### v1

@@ 2.6 Understanding the effects of fine-tuning
To develop a deeper understanding of how fine-tuning enhances performance on $\mathcal{D}_{\text{disc}}$ while damaging performance on $\mathcal{D}_{\text{cont}}$, we analyze how the prompt influences the change in loss. We find that the change in loss incurred by fine-tuning is not uniform over all prompts and depends on the likelihood that the prompt was sampled from the fine-tuning distribution $\mathcal{D}_{\text{disc}}$. In Figure 5, we see how the change in loss induced by fine-tuning varies with the likelihood of being drawn from the fine-tuning distribution. For prompts that are likely to be drawn from the fine-tuning distribution, the loss increases as we lower the likelihood. This lines up with the standard intuition that models will have stronger performance for inputs that are in-distribution and worse performance for inputs that are out-of-distribution. However, this trend does not continue forever and in fact reverses for the continuous prompts. As the likelihood continues to decrease, the model improves performance, running counter to standard intuition about out-of-distribution inputs. With this understanding of how fine-tuning affects model predictions unevenly, we can better probe what function the fine-tuned model has learned.
Figure 5: Change in loss vs density under $\mathcal{D}_{\text{disc}}$. We sample 2048 prompts of $10$ exemplars from the continuous distribution (orange) and discrete distribution (blue). For each prompt, we evaluate the log likelihood of being drawn under $\mathcal{D}_{\text{disc}}$. We also evaluate how much the loss of the $\alpha=0.5$ model changed before and after fine-tuning (scaled by the norm of the task vector). We use a binned scatterplot to show the mean and standard deviation over 10 bins of the data. We find the largest increase for the samples of the discrete distribution that are closest to the continuous distribution. We demonstrate this effect for more exemplar counts and pretrained models in Appendix B.1.
@@ 2.7 Hypothesis: Fine-tuning is suppressing solutions
where $g_{\theta}(X,y)$ is some weighting function on the discrete solution. A capability refers to whether the transformer can internally perform an algorithm of interest (i.e. discrete regression or ridge regression) and task inference refers to whether the model can correctly disambiguate which algorithm to use. If $w_{\theta}(X,y)$ were equal to the Bayes optimal $w^{*}_{\text{mix}}$ and the model perfectly implemented discrete regression and ridge regression, then $g_{\theta}(X,y)$ would be the true posterior probability that the prompt is drawn from the discrete distribution. Due to our limited mechanistic understanding of transformers, we can not test whether this is how language models compute solutions. However, we can utilize this framework as an assumption to develop insight into what function is learned by the model.
@@ 2.8 Conjugate prompting for linear regression
If the hypothesis was true, we could recover ridge regression through setting $g_{\theta}(X,y)$ to 00. Since we do not what function the transformer is precisely implementing, this is infeasible, so we try to change the prompt instead. Specifically, for $X,y$ generated under task $w$, we consider the scaled prompt $X,\gamma y$ for a scale factor $\gamma$. The scaled prompt $X,\gamma y$ is a valid linear regression problem generated under task $\gamma w$ with noise $\gamma\epsilon$. Since a sufficiently large $\gamma$ will decrease the true posterior $g(X,y)$ for all $\alpha$, we expect that $g_{\theta}(X,\gamma y)$ would be lower than $g_{\theta}(X,y)$, weighting the model output towards ridge regression. Under this scaling, the loss-optimal prediction for the scaled prompt $X,\gamma y$ would correspond to $\langle\gamma w,x_{\text{query}}\rangle$, which is simply the loss-optimal prediction for the prompt $X,y$ scaled by $\gamma$.
Figure 6: Conjugate prompting for fine-tuned models. We take transformers pretrained over $\mathcal{D}_{\text{mix}}$ for $\alpha\in\{0.2,0.5,0.8\}$ for $5000$ steps and fine-tuned over $\mathcal{D}_{\text{disc}}$ for $400$ steps. We evaluate their loss on the continuous distribution where they under-perform on ridge regression. Conjugate prompting with label scale factor $\gamma\in\{1.5,2.0\}$ recovers the pretrained solution of ridge regression, especially on lower sample counts where there is more ambiguity. We demonstrate this effect for more $\alpha$ and $\gamma$ in Appendix B.2.
Therefore, to make the model perform ridge regression instead of discrete regression, we compose our insights into the following prompting strategy. Instead of directly feeding our prompt into the model, we scale the labels $\gamma$, feed the model the scaled prompt, and scale down the model output. This should recover ridge regression if the model can perform ridge regression for the scaled prompt and if our hypothesis is true. This strategy is an instance of conjugate prompting, which we generalize in Section 3.
@@
We evaluate this strategy in our ridge regression setup in Figure 6. We find that in line with our hypothesis, conjugate prompting can help improve performance for fine-tuned models. Specifically, we observe that the strategy helps at low sample counts where it is truly ambiguous whether the prompt is drawn from the continuous or discrete distribution. At higher sample counts, the benefits of conjugate prompting are not observed since there is little ambiguity to resolve, and scaling the prompt simply tests a potentially harder task. Since we can get closer to ridge regression through conjugate prompting, we claim the ridge regression solution has not been “forgotten” but “suppressed” since it can be partially recovered through manipulating task inference.
@@ 3 Conjugate Prompting to Recover Pretraining Capabilities
In Section 2.8, we have observed that applying our model $T$ to linear regression prompts with lower likelihood of being drawn under the fine-tuning distribution yields an output with lower continuous distribution loss. We interested in generalizing this strategy to recover capabilities learnt and utilized during pretraining but not during fine-tuning. We can take advantage of this observation to design a novel prompting strategy that requires a transform $s$ from the original prompt $P$ to a new prompt $P^{\prime}$ satisfying two important properties:
@@
    (Lower likelihood) $P^{\prime}$ should have lower likelihood under the fine-tuning distribution relative to $P$. This manipulates task inference in favor of the desired solution learnt at pretraining, ensuring that when evaluating $P^{\prime}$, the model will achieve lower loss.
@@
When we “conjugate” the model by $s$, e.g. apply $s^{-1}\circ T\circ s$, we will transform the input into a space where $T$ performs the solution of interest, and then undo the original transformation, yielding a solution to the original problem that reflects the suppressed pretrained capability. Under this framework, the conjugate prompting strategy in Section 2.8 is succintly described as $s:(X,y)\to(X,\gamma y)$. When the model and capabilities of interest naturally contain such a transformation, we can design a conjugate prompting strategy which recovers pretrained capabilities.
4 Experiments on large language models
In this section, we investigate whether our understanding of fine-tuning as shifting task inference holds in large-scale language models trained on real-world data. We study two common settings for fine-tuning language models: (i) to improve their helpfulness in instruction following (Section 4.1) and (ii) to reduce their harmfulness by preventing the generation of dangerous content (Section 4.2). In each case, we show that we fine-tuning systematically seems to perform “worse” than pretraining on some tasks, but is this change catastrophic? We find that conjugate prompting of the fine-tuned model recovers some of the pretrained behavior in both cases, just like in the stylized setting of Section 2.8.
@@ 4.1 Effect of instruction tuning on in-context learning
Figure 7: Language model experiments. Left: For the in-context learning vs instruction following problem, each prompt can be solved differently by inferring the prompt as an ICL task or an IF task. Bottom: For the harmful generation problem, each harmful instruction can be solved differently by inferring the prompt as an ANSWER task (faithfully answering the problem) or REFUSE task (refusing or answering a different question).
Figure 7: Language model experiments. Left: For the in-context learning vs instruction following problem, each prompt can be solved differently by inferring the prompt as an ICL task or an IF task. Bottom: For the harmful generation problem, each harmful instruction can be solved differently by inferring the prompt as an ANSWER task (faithfully answering the problem) or REFUSE task (refusing or answering a different question).
Instruction tuning is a common fine-tuning procedure to enable pretrained LLMs to follow natural language instructions. While instruction tuning improves instruction following ability, we find that it can come at the cost of other capabilities such as in-context learning. This is particularly amplified when the two tasks are at conflict with each other. For example, suppose the prompt contains exemplars corresponding to a latent task, but the final query $x_{\text{query}}$ takes the form of an instruction (such as What is 2 + 2?). How well do models perform in-context learning in this setting?
To test this, we generate prompts using the template in Figure 7, where there are different solutions conditioned on whether the task is in-context learning (ICL) or instruction following. See Appendix C.1 for full details on how these prompts were generated and Appendix C.2 for concrete examples. We find that fine-tuned models are always less likely to perform in-context learning compared to their pre-trained counterparts: Alpaca and Vicuna-7b perform ICL on $56.75\%$ and $40.00\%$ less inputs than LLaMa-7b and OPT-IML-1.3b performs ICL on $21.00\%$ less inputs than OPT-1.3b.
We can contextualize this drop in ICL with fine-tuning under the implicit inference framework of Section 2.7. Let $\text{L}_{\text{}}(\texttt{prompt})$ denote the probability distribution over possible completions by an LLM given prompt. Let $\text{L}_{\text{IF}}$ denote this distribution conditioned on a model that always performs the task of instruction following, and let $\text{L}_{\text{ICL}}$ denote the counter part for ICL. As per our hypothesis, we can write our model $\text{L}_{\text{}}$ as
@@
where the model internally estimates $g_{\theta}$ which is the posterior likelihood of the model interpreting the latent task to be instruction following. Our hypothesis predicts that one reason instruction-tuned models are worse at ICL is because instruction-tuning increases $g_{\theta}$ for most prompts, suppressing the in-context learning capability $\text{L}_{\text{ICL}}$. Note that there might also be a change in the internal representations of $\text{L}_{\text{ICL}}$ and $\text{L}_{\text{IF}}$, but we only focus on what can be recovered by simply manipulating the task inference.
If this were true, conjugate prompting (see Section 3) can reverse the effect of $g$ and would cause the fine-tuned model to perform ICL more often.
Figure 8: Example of conjugate prompting. Left: An instruction-tuned model tends to follow the instruction instead of in-context learning. We translate the prompt to another language such as French, take the model output, and translate it to English to recover in-context learning. Right: Safety fine-tuning encourages refuse harmful instructions. We translate the prompt to another language such as Malayalam, take the model output, and translate it to English to recover harmful content generation.
Conjugate prompting to perform ICL over instruction following.
@@
We observe that the instruction tuning data for Alpaca, Vicuna, and OPT-IML are primarily in English. As a result, translating prompts to different languages satisfies the “lower likelihood” property of conjugate prompting (Section 3). Language translation also satisfies the “invertibility” property because we can simply reverse-translate the answer to English for the tasks we consider  [^2] Language translation might violate the task invertibility if we consider tasks that require specific contextual knowledge that might vary across languages. Furthermore, the pretrained models we consider are capable of performing ICL in different languages. Other than language translation, we find that the additional transformations of Leetspeak and Pig Latin (also discussed in Wei et al. (2023a)) satisfy these properties.
@@
We test whether language translation biases the fine-tuned model more towards pretrained behavior of ICL. To do so, we compute the drop in ICL frequency between the fine-tuned and pretrained counterparts across several languages, with translation implemented Google Translate (Wu et al., 2016). We present comprehensive results across 5 models, 4 non-English languages, and 2 additional transformations in in Table 1. We see that translation always results in a smaller drop in ICL frequency compared to English prompts (except for by $0.25\%$ for one model pair for one language). For example, with Alpaca, Leetspeak results in a drop of only $1.0\%$, French shows a drop of $29.00\%$, while in contrast English results in a much larger drop of $56.75\%$. This confirms that conjugate prompting can successfully revert task inference and this change can often be significant in practice. We provide a more detailed decomposition of these accuracies in Appendix C.3.
@@
Table 1: Measuring in-context learning vs instruction following. We report the accuracy that the model provides the correct first word completion to the in-context learning task. Accuracies are taken over 400 samples from the 4 combinations of 2 ICL tasks and 2 instruction following tasks. We find that over model pairs, instruction-tuned models are more likely to perform the instruction-following task than the pretrained models. Fine-tuned models in English are always least likely to perform the in-context learning task and exhibit the largest drop in likelihood to perform the in-context task in relation to their pre-trained counterpart (except for by 0.25% compared to LLaMa/Vicuna in Dutch).
Pretrained Model Fine-tuned Model Language Pretrained ICL Accuracy Fine-tuned ICL Accuracy Change in ICL Task
@@
4.2 Effects of Reinforcement Learning from Human Feedback On Harmful Content Generation
Since models are pretrained to model any harmful text found on the internet, they are not immediately safe for deployment. As such, models are typically fine-tuned to produce text which reflect human preferences through Reinforcement Learning from Human Feedback (RLHF).
Does this fit within our framework? Consider a prompt like “How do I build a bomb?”. If we refer to $\text{L}_{\text{ANSWER}}$ as the capability that attempts to answer the question while $\text{L}_{\text{REFUSE}}$ is the solution that refuses to answer the question, we can idealize the model’s completion as

### v2

@@ 2.6 Understanding the effects of fine-tuning
To develop a deeper understanding of how fine-tuning enhances performance on $\mathcal{D}_{\text{disc}}$ while damaging performance on $\mathcal{D}_{\text{cont}}$, we analyze how the prompt influences the change in loss. We find that the change in loss incurred by fine-tuning is not uniform and depends on the likelihood that the prompt was sampled from the fine-tuning distribution $\mathcal{D}_{\text{disc}}$. In Figure 5, we see how the change in loss induced by fine-tuning varies with the likelihood of being drawn from the fine-tuning distribution. For prompts that are likely to be drawn from the fine-tuning distribution, the loss increases as we lower the likelihood. This lines up with the standard intuition that models will have stronger performance for inputs that are in-distribution and worse performance for inputs that are out-of-distribution. However, this trend does not continue forever and in fact reverses for the continuous prompts. As the likelihood continues to decrease, the model improves performance, running counter to standard intuition about out-of-distribution inputs. With this understanding of how fine-tuning affects model predictions unevenly, we can better probe what function the fine-tuned model has learned.
Figure 5: Change in loss vs density under $\mathcal{D}_{\text{disc}}$. We sample 2048 prompts of $10$ exemplars from $\mathcal{D}_{\text{cont}}$ and $\mathcal{D}_{\text{disc}}$ (blue) and evaluate the log likelihood of being drawn from $\mathcal{D}_{\text{disc}}$. We also evaluate how much the loss of the $\alpha=0.5$ model changed before and after fine-tuning (scaled by the norm of the task vector). The binned scatterplot shows the mean and standard deviation for 10 bins; the largest increase is for $\mathcal{D}_{\text{disc}}$ samples closest to $\mathcal{D}_{\text{cont}}$. More examples in Appendix C.1.
@@ 2.7 Hypothesis: Fine-tuning is suppressing solutions
where $g_{\theta}(X,y)$ is some weighting function on the discrete solution estimating the posterior probability that the prompt is drawn from $\mathcal{D}_{\text{disc}}$. A capability refers to whether the transformer can internally perform an algorithm of interest (i.e. discrete regression or ridge regression) and task inference refers to whether the model can correctly disambiguate which algorithm to use. Due to limited mechanistic understanding, we can not test whether this is how language models compute solutions. However, we can utilize this as an assumption to understand what is learned by the model.
@@ 2.8 Conjugate prompting for linear regression
If the hypothesis was true, we could recover ridge regression through setting $g_{\theta}(X,y)$ to 00. Since we do not know what function the transformer is precisely implementing, this is infeasible, so we try to change the prompt instead. Specifically, for $X,y$ generated under task $w$, we consider the scaled prompt $X,\gamma y$ for a scale factor $\gamma$. The scaled prompt $X,\gamma y$ is a valid regression problem generated under task $\gamma w$ with noise $\gamma\epsilon$. Since a sufficiently large $\gamma$ will decrease the true posterior $g(X,y)$ for all $\alpha$, we expect that $g_{\theta}(X,\gamma y)$ would be lower than $g_{\theta}(X,y)$, weighting the output towards ridge regression. Under this scaling, the loss-optimal prediction for the scaled prompt $X,\gamma y$ would correspond to $\langle\gamma w,x_{\text{query}}\rangle$, which is the loss-optimal prediction for the prompt $X,y$ scaled by $\gamma$.
Refer to caption Refer to caption Refer to caption
Figure 6: Conjugate prompting for regression. We take transformers pretrained over $\mathcal{D}_{\text{mix}}$ for $\alpha\in\{0.2,0.5,0.8\}$ for $5000$ steps and fine-tuned over $\mathcal{D}_{\text{disc}}$ for $400$ steps. We evaluate their loss on the continuous distribution where they under-perform ridge regression. Conjugate prompting with scale factor $\gamma\in\{1.5,2.0\}$ recovers the pretrained solution of ridge regression, especially on lower sample counts with more ambiguity. We demonstrate this effect for more $\alpha,\gamma$ in Appendix C.2.
Therefore, to make the model perform ridge regression, we compose our insights into the following prompting strategy. Instead of directly feeding our prompt into the model, we scale the labels $\gamma$, feed the model the scaled prompt, and scale down the model output. This should recover ridge regression if the model can perform ridge regression for the scaled prompt and if our hypothesis is true. This strategy is an instance of conjugate prompting, which we generalize in Section 3.
@@
We evaluate conjugate prompting Figure 6. In line with our hypothesis, conjugate prompting can help improve performance for fine-tuned models. Specifically, we observe that the strategy helps at low sample counts where the fine-tuned model is more uncertain if the prompt is from the continuous or discrete distribution. We suspect that at higher sample counts, the fine-tuned model has better inferred the task and conjugate prompting simply tests a harder task. Since we can get closer to ridge regression through conjugate prompting, we claim the ridge regression solution has not been “forgotten” but “suppressed” since it can be partially recovered through manipulating task inference.
@@ 3 Conjugate Prompting to Recover Pretraining Capabilities
In Section 2.8, we observed that applying our model $T$ to regression prompts with lower likelihood under the fine-tuning distribution yields lower continuous distribution loss. We are interested in generalizing this to recover pretraining capabilities not utilized after fine-tuning. We design a prompting strategy that uses a transform $s$ from prompt $P$ to a new prompt $P^{\prime}$ satisfying two properties:
@@
    (Lower likelihood) $P^{\prime}$ should have lower likelihood under the fine-tuning distribution to shift task inference in favor of the pretraining solution.
@@
When we “conjugate” the model by $s$ (apply $s^{-1}\circ T\circ s$), we transform the input into a space where $T$ performs the solution of interest and then undo the original transformation, yielding a solution that reflects the suppressed pretrained capability. The conjugate prompting strategy in Section 2.8 is succintly described as $s:(X,y)\to(X,\gamma y)$. When the model and training distributions naturally contain such a transformation, we can utilize conjugate prompting to recover pretrained capabilities.
4 Experiments on language models
In this section, we investigate whether our understanding of fine-tuning as shifting task inference holds in large-scale language models trained on real-world data. We study three common settings for fine-tuning language models: (i) improving helpfulness in instruction following (Section 4.1) (ii) improving coding capabilities with domain fine-tuning (Section 4.2) and (ii) reducing harmfulness by preventing the generation of dangerous content (Section 4.3). In each case, though fine-tuning seems to perform “worse” than pretraining on some tasks, conjugate prompting can recover some of the pretrained behavior from the fine-tuned model just like the stylized setting of Section 2.8.
@@ 4.1 Effect of instruction tuning on in-context learning
Figure 7: Language model experiments. We test forgetting of in-context learning after instruction tuning (left), natural language reasoning after code fine-tuning (middle), and toxic generation after safety fine-tuning (right). For each, the blue reflects the fine-tuning task while the orange reflects the forgotten task.
Instruction tuning is a common procedure to enable pretrained LLMs to follow natural language instructions. While instruction tuning improves instruction following (IF) ability, we find that it can come at the cost of other capabilities such as in-context learning. This is particularly amplified when the two tasks are at conflict with each other. For example, suppose the prompt contains exemplars corresponding to a latent task, but the final query $x_{\text{query}}$ takes the form of an instruction (such as What is 2 + 2?). How well do models perform in-context learning in this setting?
To test this, we consider a controlled experiment where prompts follow the template in Figure 7 with different solutions if the task is in-context learning (ICL) vs instruction following (IF) (evaluation details in Appendix D.1, examples in Appendix D.2). We find that fine-tuned models are always less likely to perform in-context learning compared to their pre-trained counterparts: Alpaca (Taori et al., 2023) and Vicuna-7b (Chiang et al., 2023) perform ICL on $56.75\%$ and $40.00\%$ less inputs than LLaMa-7b (Touvron et al., 2023a) and OPT-IML-1.3b (Iyer et al., 2023) performs ICL on $21.00\%$ less inputs than OPT-1.3b (Zhang et al., 2022).
We can contextualize this drop in ICL with fine-tuning under the implicit inference framework of Section 2.7. Let $\text{L}_{\text{}}(\texttt{prompt})$ denote the distribution over possible completions by a model $\text{L}_{\text{}}$ given prompt. Let $\text{L}_{\text{IF}}$ denote this distribution conditioned on a model that always follows instructions, and $\text{L}_{\text{ICL}}$ be the same for ICL. As per our hypothesis, we can write our model $\text{L}_{\text{}}$ as
@@
where the model internally estimates $g_{\theta}$ which is the posterior likelihood of the model interpreting the latent task to be instruction following. Our hypothesis predicts that one reason instruction-tuned models are worse at ICL is because instruction-tuning increases $g_{\theta}$ for most prompts, suppressing the in-context learning capability $\text{L}_{\text{ICL}}$. Note that there might also be a change in the internal representations of $\text{L}_{\text{ICL}}$ and $\text{L}_{\text{IF}}$, but we only focus on what can be recovered by simply manipulating the task inference. If our hypothesis holds, conjugate prompting (see Section 3) can reverse the effect of $g$ and would cause the fine-tuned model to perform ICL more often.
Conjugate prompting to perform ICL.
@@
We observe that the instruction tuning data for Alpaca, Vicuna, and OPT-IML is primarily in English. Therefore, translating to different languages satisfies the “lower likelihood” property as well as the “invertibility” property of conjugate prompting because we can simply translate the answer to English  [^2] Translation can violate invertibility for tasks that require contextual knowledge that varies across languages. Other than language translation, we consider the additional transformations of Leetspeak and Pig Latin (discussed in Wei et al. (2023a)).
@@
We test whether language translation can recover the pretrained behavior of ICL. To do so, we compute the drop in ICL frequency between the English fine-tuned and pretrained counterparts across 5 models prompted under 4 non-English languages and 2 additional transformations in in Table 1 (translation implemented with Google Translate (Wu et al., 2016)). We see that translation almost always results in a smaller drop in ICL frequency compared to English prompts. For example, with Alpaca, Leetspeak results in a drop of only $1.0\%$, French results in a drop of $29.00\%$, while English results in a drop of $56.75\%$. This confirms that conjugate prompting can successfully shift task inference in practice. We provide a more detailed decomposition in Appendix D.3.
@@
Table 1: Measuring in-context learning vs instruction following. We report the accuracy of first word completion for in-context learning task. Accuracies are taken over 400 samples and 4 ICL vs IF tasks. Instruction-tuned models are least likely to perform in-context learning task when prompted in English (except for by 0.25% compared to Vicuna in Dutch) and almost always exhibit the largest drop in likelihood to perform the ICL task.
Pretrained Fine-tuned Language Pretrained ICL Acc Fine-tuned ICL Acc Drop in ICL Task
@@
Figure 8: Example of conjugate prompting. We highlight how conjugate prompting operates for circumventing forgetting from code fine-tuning where we find that the fine-tuned model may not perform the reasoning task at hand. Instead of directly prompting the model, we first translate the input to a different language (such as Spanish), apply the model, and translate the output back.
4.2 Effects of code fine-tuning on natural language reasoning
To demonstrate a more natural instance of catastrophic forgetting, we consider what happens to a language model after we fine-tune on code. If we refer to $\text{L}_{\text{REASON}}$ as the capability that solves a natural language reasoning task while $\text{L}_{\text{CODE}}$ does the same for coding, we can idealize the model’s completion as
$\text{L}_{\text{}}(\texttt{prompt})=g_{\theta}(\texttt{prompt})\text{L}_{\text{CODE}}(\texttt{prompt})+(1-g_{\theta}(\texttt{prompt}))\text{L}_{\text{REASON}}(\texttt{prompt})$
Conjugate prompting for MNLI.
To test forgetting, we use the XNLI benchmark (Conneau et al., 2018), a multi-lingual version of MNLI (Williams et al., 2018) from GLUE (Wang et al., 2019) that tests sentence entailment (evaluation details in Appendix E.1, examples in Appendix E.2).
When we compare LLaMa-2 (Touvron et al., 2023b) against its English code fine-tuned variant Code LLaMa (Rozière et al., 2023), the model gets lower performance on English prompts, performing $8.36\%$ worse (Table 2). However, for French, Spanish, and German inputs, the accuracy changes by less than $2\%$ in magnitude. In fact, the accuracy of Code LLaMa on Spanish and French XNLI slightly increases after fine-tuning, possibly from increased reasoning capabilities associated with code training (Fu & Khot, 2022; Ma et al., 2023) combined with better task inference. For this reasoning task, it is preferable to prompt the fine-tuned model in Spanish instead of English.
4.3 Effects of RLHF On Harmful Content Generation
Since models are pretrained on harmful text found on the internet, they are typically fine-tuned to reflect human preferences through RLHF. Does this fit within our framework? If we refer to $\text{L}_{\text{ANSWER}}$ as the capability that attempts to answer an instruction while $\text{L}_{\text{REFUSE}}$ is the solution that refuses to answer the question, we can idealize the model’s completion as below. Safety fine-tuning induces a drop in answering and can be studied under our framework similarly to forgetting.